# Clock Auctions Augmented with Unreliable Advice

### v1

@@ 2 Preliminaries
We design clock auctions for this problem. A clock auction is a dynamic mechanism proceeding over several “rounds”. In each round $t$, each bidder $i$ is offered a personalized “clock” price $p_{i,t}$ which represents the current amount they would need to pay if the auction were to terminate now. At the outset of the auction, the clock price $p_{i,0}$ for each bidder $i$ is initialized to some arbitrarily small positive value $v_{\min}$ and each bidder is placed in an “active” bidder set $A$. The clock prices are non-decreasing throughout the auction, i.e., $p_{i,t}\geq p_{i,t-1}$ for all rounds $t$ [^1] We assume that each bidder with positive value has value above a publicly known minimum value $v_{\min}$, which is consistent with [BLP2009] and [CGS22]. As such, rather than continuous prices, we may increment prices by an arbitrarily small positive value $\delta$, say, $v_{\min}/n^{2}$. As argued in [CGS22], discretizing prices to this range causes negligible loss in approximation since the optimal social welfare is lower bounded by $v_{\min}$. In each round $t$, the auctioneer announces a price $p_{i,t}$ for each bidder $i\in A$ and bidder $i$ can then choose to remain in the auction (staying “active”) or permanently exit (becoming “inactive” and being removed from $A$). If $i$ exits at any point in the auction she will not receive the service and pays nothing. The auction then terminates when $A\in\mathcal{F}$, i.e., all active bidders may be feasibly served. Note that in each round the clock price for each bidder can only depend on public information, i.e., the feasibility constraint, the history of prices, and the points at which bidders exited the auction. We use $\text{rev}(S,\mathbf{p_{t}})=\sum_{i\in S\cap A}{p_{i,t}}$ to denote the revenue of a set $S$ at current price vector $\mathbf{p_{t}}$. In other words, the revenue of a set $S$ in round $t$ of the auction is the sum of the clock prices in round $t$ of the active active bidders in $S$.
To analyze the performance of a clock auction $\mathcal{M}$, we consider its worst-case approximation guarantee. On an instance $I$, let $I$ denote the welfare obtained by the clock auction on $I$ and $\texttt{OPT}(I)$ denote the maximum social welfare among all feasible sets in $I$. Then, $\mathcal{M}$ obtains $\alpha$-approximation of the optimal social welfare for a class of instances $\mathcal{I}$ if
@@
The water-filling clock auction
The central result of [CGS22] is a deterministic clock auction called the “water-filling clock auction”, i.e., WFCA, which achieves the (near) optimal approximation guarantee of $O(\log{n})$ any any family of instances where $\mathcal{F}$ is downward-closed. The WFCA computes the revenue of each set in each round and marks the set of bidders $W$ with highest revenue as the “conditional winners” and increments the price of the lowest priced bidders outside $W$. This, intuitively, makes sense since the revenue of a set is a lower bound on its remaining possible welfare. As the WFCA is an important subroutine in our auctions, we include pseudocode for it here as Subroutine 1 for completion.
@@ 10                $A\leftarrow A\setminus\{i\}$
We summarize two critical features of the WFCA shown in [CGS22] that we use in our analysis as Lemma 1 below. We note that the authors of [CGS22] show that the WFCA achieves a $4\log{n}$-approximation [CGS22, Theorem 2]. The first theorem demonstrates a slightly stronger guarantee that is more useful for our bounds, namely that the welfare achieved is a $2H_{n}$-approximation where $H_{n}=\sum_{i=1}^{n}{1/i}$ is the $n$-th harmonic number. It is known that $H_{n}$ is asymptotically $O(\log{n})$.
Lemma 1 (CGS22).
@@
The learning-augmented framework
In this work we adopt the learning-augmented framework and study clock-auctions that are also equipped with a (potentially erroneous) prediction $\hat{\texttt{OPT}}$ regarding the highest welfare feasible set to serve. Given the predictions and an instance, we denote the welfare achieved by a mechanism $\mathcal{M}$ as $\mathcal{M}(I,\hat{\texttt{OPT}})$ and we evaluate the performance of $\mathcal{M}$ using its robustness and consistency. [^2] We note that the predictions used by mechanisms are public, i.e., the mechanism designer and bidders observe the predictions. Although our clock auctions utilize the predictions to guide the price-increase process, the bidders still face only monotonically increasing clock prices and, thus, they still have a simple interface with an obviously dominant strategy of exiting the auction when the clock price becomes undesirable (higher than their private value).
@@
The consistency of a mechanism refers to the worst-case approximation ratio of the mechanism when the prediction that it is provided with is accurate, i.e., $=\texttt{OPT}$ OPT. Mathematically,
$\text{consistency}(\mathcal{M})=\max_{I}\frac{v(\texttt{OPT}(I))}{v(\mathcal{M}(I,\texttt{OPT}(I)))}.$ (1)
Additionally, we consider a stronger notion of consistency first proposed by [MNS07], which we denote as consistency. It is defined to be the worst-case the approximation ratio with respect to the value of the predicted set, regardless of the correctness of the prediction. Mathematically,
@@ 3 Clock Auctions Achieving the Best of Both Worlds
In this section, we present a deterministic clock auction augmented with a prediction $\hat{\texttt{OPT}}\in\mathcal{F}$ suggesting what is the social welfare-maximizing feasible set. Our auction leverages this prediction to achieve a good trade-off between consistency (see Equation (1)) and robustness. Recall that without predictions, no mechanism can achieve an $O(\log^{1-\epsilon}n)$-approximation for any arbitrarily small $\epsilon$. The hope is to bypass this logarithmic barrier whenever the prediction is accurate, without hurting the robustness too much. One naive approach is to just serve the predicted set, which would yield 1-consistency, but it is not hard to see that it would suffer from unbounded robustness. On the other hand, ignoring the prediction and just running the WFCA would give us $O(\log n)$ consistency with $O(\log n)$ robustness. A guarantee one can hope for is small constant consistency with $O(\log n)$ robustness, which is also referred to as "best-of-both-worlds." This is what our clock auction, formally defined in Mechanism 2, achieves. In fact, we can let the consistency be arbitrarily $1+\epsilon$ for an arbitrarily small constant $\epsilon>0$, while maintaining $O(\frac{1}{\epsilon}\log n)$-robustness. We also complement our positive result with a lower bound showing that this robustness bound is tight with respect to its dependence on both $n$ and $\epsilon$.
@@
For any downward-closed set system, Mechanism 2 with parameter $\gamma=\frac{10(1+\epsilon)}{9\epsilon}$ is $(1+\epsilon)$-consistent and $O(\frac{1}{\epsilon}\log{n})$-robust for any constant $\epsilon>0$.
The mechanism, named FollowTheUnpredictedLeader, is a clock auction that alternates between raising the price for bidders in the predicted set and the unpredicted set uniformly, while maintaining a balanced ratio between the target revenue from the predicted set and the rejected welfare of the unpredicted set. At a high level, in every iteration, the mechanism first pushes the unpredicted bidders only until we are about to lose the robustness guarantees based on the revenue already on hand. It then further pushes the unpredicted bidders, aiming to ensure the robustness guarantees is satisfied by the unpredicted set solely. If the unpredicted bidders all reject the offers during the process, the mechanism outputs the remaining active bidders in the predicted set. Otherwise, the mechanism switches to pushing the predicted set to match the current revenue obtained. If such revenue is achieved, we repeat the process; otherwise (if all predicted bidders reject), we run the WFCA on the remaining unpredicted bidders. We implicitly assume that the bidders choose the obviously dominant strategy to exit the auction whenever the clock price exceeds their value, and we learn the values at the exit of the bidders. Please see Mechanism 2 for a formal description.
@@ 6      $R_{t}\leftarrow 10\cdot R_{t-1}$
     Run $\textsc{UniformPrice}(N\setminus\hat{\texttt{OPT}},\mathbf{p})$ until $\max_{F}v(F\setminus N)\in[R_{t}\gamma H_{n},2R_{t}\gamma H_{n})$333we can make sure the rejected welfare are within this range by stopping when current price of active unpredicted bidders reaches $R_{t}\gamma H_{n}$
@@ 6                $S\leftarrow S\setminus\{i\}$
We first show that Line 2 does not affect neither consistency nor robustness by too much.
@@
We first note that the social welfare of the predicted set is the same before and after completing the transformation in line 5 of Mechanism 2. Then if the predicted set was the optimal set before the transformation then the value of the optimal set does not decrease. So suppose some non-predicted set $O^{*}$ is optimal and denote its total welfare $v(O^{*})$. If at least half of the welfare contained in $O^{*}$ comes from bidders contained in the predicted set then after the transformation the predicted set has welfare at least $v(O^{*})/2$ and, hence, the optimal set has welfare at least $v(O^{*})/2$ after the transformation. On the other hand, if more than half of the welfare contained in $O^{*}$ comes from bidders not in the predicted set then after the transformation $O^{*}$ still has social welfare at least $v(O^{*})/2$. ∎
@@
Robustness Argument
We first establish the robustness of Mechanism 2. To begin, we demonstrate the robustness guarantee if a subset of the predicted set is outputted (Lemma 3). For cases where an unpredicted set is outputted, we argue that the revenue obtained at the time point where $\hat{\texttt{OPT}}=\emptyset$ provides a good approximation of the best welfare achievable from any rejected subset (Lemma 4). We then utilize the revenue monotonicity of WFCA to achieve the desired robustness bound (Lemma 5).
@@
In the case that Mechanism 2 outputs a subset of unpredicted bidders it must arrive at line 2. As such, we can decompose the social welfare from the optimal (unpredicted) set into two components, the portion of the welfare coming from bidders rejected before Mechanism 2 runs the Water-filling Clock Auction and the portion of the welfare coming from bidders who participate in the WFCA. From Lemma 4 we know that the revenue reached by Mechanism 2 before running the WFCA is within a $\frac{9}{10(2\gamma+1)H_{n}}$-factor of the welfare rejected from the optimal set before running the WFCA. Moreover, from Lemma 1 we have that the revenue reached by Mechanism 2 after running the WFCA is weakly higher than the revenue reached before running the WFCA. As such, since the welfare obtained by serving a set of bidders is always weakly higher than the revenue collected from these bidders we have that the social welfare obtained by Mechanism 2 is within a $\frac{9}{10(2\gamma+1)H_{n}}$-factor of the welfare rejected from the optimal set before running the WFCA. Finally, we have from Lemma 1 that the social welfare achieved by running the WFCA only on bidders who are active when line 2 is reached is within a $1/(2H_{n})$-factor of the optimal social welfare achievable from these bidders. Combining these two guarantees, we have that the social welfare obtained by Mechanism 2 is within a $\frac{9}{10(2\gamma+1)H_{n}}+\frac{1}{2H_{n}}$-factor of the optimal social welfare. ∎
Consistency Argument
We now establish the consistency of Mechanism 2. The high-level idea is to upper bound the total rejected welfare when the mechanism terminates and argue that: (i) If the predicted set is indeed optimal, the mechanism will always output a subset of the predicted set. (ii) The revenue, which serves as a lower bound of welfare, from the output set is sufficient to guarantee the desired consistency.
@@
$\frac{10^{k-1}\cdot\gamma H_{n}-10^{k-1}\cdot\frac{10}{9}H_{n}}{10^{k-1}\cdot\gamma H_{n}}=\frac{\gamma-\frac{10}{9}}{\gamma}.\qed$
@@
Proof of Theorem 1.
First observe that $\epsilon>0$ ensures that $\gamma>10/9$. In the case that the predicted set, indeed, optimal, Lemma 7 gives that some subset of the predicted set is served. Moreover, we obtain that the total social welfare of served bidders is within a factor
@@
of the total social welfare in the predicted set, thereby guaranteeing $(1+\epsilon)$-consistency. On the other hand, if some subset of the predicted set is served, Lemma 3 guarantees that the total social welfare obtained by Mechanism 2 is within a factor $\frac{9}{100(2\gamma+1)H_{n}}$ of the optimal social welfare. Since $H_{n}=O(\log{n})$ and $\gamma$ is a constant for any choice of constant $\epsilon$ we obtain $O(\log{n})$-robustness. Finally, if some subset of an unpredicted set is served, Lemma 5 guarantees that Mechanism 2 obtains welfare within a factor $\frac{9}{10(2\gamma+1)H_{n}}+\frac{1}{2H_{n}}$ of the optimal social welfare. Again, since $H_{n}=O(\log{n})$ and $\gamma$ is constant for any constant $\epsilon$ we obtain $O(\log{n})$-robustness in this case, completing the proof. ∎

### v2

@@ 2 Preliminaries
We design clock auctions for this problem. A clock auction is a dynamic mechanism proceeding over several rounds. In each round $t$, each bidder $i$ is offered a personalized “clock” price $p_{i,t}$, which represents the current amount they would need to pay if the auction were to terminate at that moment. At the outset of the auction, the clock price $p_{i,0}$ for each bidder $i$ is initialized to some arbitrarily small positive value $v_{\min}$, and each bidder is placed in an “active” bidder set $A$. The clock prices are non-decreasing throughout the auction, i.e., $p_{i,t}\geq p_{i,t-1}$ for all rounds $t$ [^1] We assume that each bidder with positive value has a value above a publicly known minimum value $v_{\min}$, which is consistent with [BLP2009] and [CGS22]. Thus, rather than continuous prices, we may increment prices by an arbitrarily small positive value $\delta$, say, $v_{\min}/n^{2}$. As argued in [CGS22], discretizing prices to this range causes negligible loss in approximation since the optimal social welfare is lower-bounded by $v_{\min}$. In each round $t$, the auctioneer announces a price $p_{i,t}$ for each bidder $i\in A$, and bidder $i$ can then choose to remain in the auction (staying “active”) or permanently exit (becoming “inactive” and being removed from $A$). If $i$ exits at any point in the auction, they will not receive the service and pay nothing. The auction then terminates when $A\in\mathcal{F}$, i.e., all active bidders may be feasibly served.
Note that in each round, the clock price for each bidder can only depend on public information, i.e., the feasibility constraint, the history of prices, and the points at which bidders exited the auction. We use $\text{rev}(S,\mathbf{p_{t}})=\sum_{i\in S\cap A}{p_{i,t}}$ to denote the revenue of a set $S$ at the current price vector $\mathbf{p_{t}}$. In other words, the revenue of a set $S$ in round $t$ of the auction is the sum of the clock prices in round $t$ of the active bidders in $S$.
To analyze the performance of a clock auction $\mathcal{M}$, we consider its worst-case approximation guarantee. On an instance $I$, let $I$ denote the welfare obtained by the clock auction on $I$ and $\texttt{OPT}(I)$ denote the maximum social welfare among all feasible sets in $I$. Then, $\mathcal{M}$ obtains an $\alpha$-approximation of the optimal social welfare for a class of instances $\mathcal{I}$ if
@@
The Water-Filling Clock Auction.
The central result of [CGS22] is a deterministic clock auction called the “Water-Filling Clock Auction” (WFCA), which achieves the (near) optimal approximation guarantee of $O(\log{n})$ for any family of instances where $\mathcal{F}$ is downward-closed. The WFCA computes the revenue of each set in each round and marks the set of bidders $W$ with the highest revenue as the “conditional winners” and increments the price of the lowest-priced bidders outside $W$. This, intuitively, makes sense since the revenue of a set is a lower bound on its remaining possible welfare. As the WFCA is an important subroutine in our auctions, we include pseudocode for it here as Subroutine 1 for completion.
@@ 10                $A\leftarrow A\setminus\{i\}$
We summarize two critical features of the WFCA shown in [CGS22] that we use in our analysis as Lemma 1 below. We note that the authors of [CGS22] show that the WFCA achieves a $4\log{n}$-approximation [CGS22, Theorem 2]. The proof of this theorem demonstrates a slightly stronger guarantee that is more useful for our bounds, namely that the welfare achieved is a $2H_{n}$-approximation, where $H_{n}=\sum_{i=1}^{n}{1/i}$ is the $n$-th harmonic number. It is known that $H_{n}$ is asymptotically $O(\log{n})$.
Lemma 1 ([CGS22]).
@@
The Learning-Augmented Framework.
In this work, we adopt the learning-augmented framework and study clock auctions that are also equipped with a (potentially erroneous) prediction $\hat{\texttt{OPT}}$ regarding the feasible set of highest welfare. Given the predictions and an instance, we denote the welfare achieved by a mechanism $\mathcal{M}$ as $\mathcal{M}(I,\hat{\texttt{OPT}})$, and we evaluate the performance of $\mathcal{M}$ using its robustness and consistency. [^2] We note that the predictions used by mechanisms are public, i.e., the mechanism designer and bidders observe the predictions. Although our clock auctions utilize the predictions to guide the price-increase process, the bidders still face only monotonically increasing clock prices and, thus, they still have a simple interface with an obviously dominant strategy of exiting the auction when the clock price becomes undesirable (higher than their private value).
@@
The consistency of a mechanism refers to the worst-case approximation ratio of the mechanism when the prediction it is provided with is accurate, i.e., $=\texttt{OPT}$ OPT. Mathematically,
$\text{consistency}(\mathcal{M})=\max_{I}\frac{v(\texttt{OPT}(I))}{\mathcal{M}(I,\texttt{OPT}(I))}.$ (1)
Additionally, we consider a stronger notion of consistency first proposed by [MNS07], which we denote as consistency. It is defined to be the worst-case approximation ratio with respect to the value of the predicted set, regardless of the correctness of the prediction. Mathematically,
@@ 3 Clock Auctions Achieving the Best of Both Worlds
In this section, we present a deterministic clock auction augmented with a prediction $\hat{\texttt{OPT}}\in\mathcal{F}$ that suggests the feasible set which maximizes social welfare. Our auction leverages this prediction to achieve a favorable trade-off between consistency (see Equation (1)) and robustness. Recall that without predictions, no mechanism can achieve an $O(\log^{\tau}n)$-approximation for any constant $\tau<1$. The aim is to bypass this logarithmic barrier whenever the prediction is accurate, without significantly sacrificing robustness.
One naive approach is to simply serve the predicted set, which would yield 1-consistency; however, this would suffer from unbounded robustness. Conversely, ignoring the prediction and simply running the WFCA would provide $O(\log n)$ consistency and $O(\log n)$ robustness. A desirable outcome is small constant consistency with $O(\log n)$ robustness, often referred to as a “best-of-both-worlds” guarantee. This is precisely what our clock auction, formally defined in Mechanism 2, achieves. In fact, we can set the consistency arbitrarily close to $1+\epsilon$ for any small constant $\epsilon>0$, while maintaining $O\left(\frac{1}{\epsilon}\log n\right)$-robustness. We complement this positive result with a lower bound, showing that this robustness bound is tight with respect to its dependence on both $n$ and $\epsilon$.
@@
For any downward-closed set system, Mechanism 2 with parameter $\gamma=\frac{10(1+\epsilon)}{9\epsilon}$ is $(1+\epsilon)$-consistent and $O\left(\frac{1}{\epsilon}\log{n}\right)$-robust for any constant $\epsilon>0$.
The mechanism, called FollowTheUnpredictedLeader, is a clock auction that alternates between raising the price for bidders in the predicted and unpredicted sets, while maintaining a balanced ratio between the target revenue from the predicted set and the rejected welfare of the unpredicted set. At a high level, in each iteration, the mechanism first sets a new, higher revenue “target” $R_{t}$ and pushes the unpredicted bidders only until the robustness guarantee based on current revenue is nearly at risk. It then further pushes the unpredicted bidders, aiming to ensure the robustness guarantee is satisfied solely by the unpredicted set. If all unpredicted bidders reject the offers during this process, the mechanism outputs the remaining active bidders in the predicted set. Otherwise, it switches to pushing the predicted set to match the current revenue obtained. If the target revenue is met, the process repeats; if all predicted bidders reject, the WFCA is run on the remaining unpredicted bidders. We assume that bidders choose the obviously dominant strategy of exiting the auction whenever the clock price exceeds their value, and we learn each bidder’s value upon their exit.
For a formal description, please refer to Mechanism 2.
@@ 6      $R_{t}\leftarrow 10\cdot R_{t-1}$
     Run $\textsc{UniformPrice}(N\setminus\hat{\texttt{OPT}},\mathbf{p})$ until $\max_{F}v(F\setminus N)\in[R_{t}\gamma H_{n},2R_{t}\gamma H_{n})$
       // Reject ‘‘safe’’ amount of welfare to determine minimum total value in optimal set [^3] We can make sure the total welfare in some unpredicted set is at least $R_{t}\gamma H_{n}$ but the rejected welfare is no more than $2R_{t}\gamma H_{n}$ by stopping when current price of active unpredicted bidders reaches $R_{t}\gamma H_{n}$.
@@ 6                $S\leftarrow S\setminus\{i\}$
We first show that Line 2 affects neither consistency nor robustness by too much.
@@
We first note that the social welfare of the predicted set is the same before and after completing the transformation in line 5 of Mechanism 2. Then if the predicted set was the optimal set before the transformation, then the value of the optimal set does not decrease. So suppose some non-predicted set $O^{*}$ is optimal and denote its total welfare $v(O^{*})$. If at least half of the welfare contained in $O^{*}$ comes from bidders contained in the predicted set then after the transformation the predicted set has welfare at least $v(O^{*})/2$ and, hence, the optimal set has welfare at least $v(O^{*})/2$ after the transformation. On the other hand, if more than half of the welfare contained in $O^{*}$ comes from bidders not in the predicted set then after the transformation $O^{*}$ still has social welfare at least $v(O^{*})/2$. ∎
@@
Robustness Argument.
We first establish the robustness of Mechanism 2. To begin, we demonstrate the robustness guarantee if a subset of the predicted set is output (Lemma 3). For cases where an unpredicted set is output, we argue that the revenue obtained at the time point where $\hat{\texttt{OPT}}=\emptyset$ provides a good approximation of the best welfare achievable from any rejected subset (Lemma 4). We then utilize the revenue monotonicity of WFCA to achieve the desired robustness bound (Lemma 5).
@@
In case Mechanism 2 outputs a subset of unpredicted bidders it must arrive at line 2. As such, we can decompose the social welfare from the optimal (unpredicted) set into two components, the portion of the welfare coming from bidders rejected before Mechanism 2 runs the Water-filling Clock Auction and the portion of the welfare coming from bidders who participate in the WFCA. From Lemma 4 we know that the revenue reached by Mechanism 2 before running the WFCA is within a $\frac{9}{10(2\gamma+1)H_{n}}$-factor of the welfare rejected from the optimal set before running the WFCA. Moreover, from Lemma 1 we have that the revenue reached by Mechanism 2 after running the WFCA is weakly higher than the revenue reached before running the WFCA. As such, since the welfare obtained by serving a set of bidders is always weakly higher than the revenue collected from these bidders we have that the social welfare obtained by Mechanism 2 is within a $\frac{9}{10(2\gamma+1)H_{n}}$-factor of the welfare rejected from the optimal set before running the WFCA. Finally, we have from Lemma 1 that the social welfare achieved by running the WFCA only on bidders who are active when line 2 is reached is within a $1/(2H_{n})$-factor of the optimal social welfare achievable from these bidders. Combining these two guarantees, we have that the social welfare obtained by Mechanism 2 is within a $\frac{9}{10(2\gamma+1)H_{n}}+\frac{1}{2H_{n}}$-factor of the optimal social welfare. ∎
Consistency Argument.
We now establish the consistency of Mechanism 2. The high-level idea is to upper bound the total rejected welfare when the mechanism terminates and argue that: (i) If the predicted set is indeed optimal, the mechanism will always output a subset of the predicted set; and (ii) the remaining welfare in the predicted set (when it is output) is high as a result of us “favoring” the predicted set in Mechanism 2.
@@
$\frac{10^{k-1}\cdot\gamma H_{n}-10^{k-1}\cdot\frac{10}{9}H_{n}}{10^{k-1}\cdot\gamma H_{n}}=\frac{\gamma-\frac{10}{9}}{\gamma},$
which completes the proof. ∎
@@
Proof.
(Proof for Theorem 1) First observe that $\epsilon>0$ ensures that $\gamma>10/9$. In the case that the predicted set, indeed, optimal, Lemma 7 gives that some subset of the predicted set is served. Moreover, we obtain that the total social welfare of served bidders is within a factor
@@
of the total social welfare in the predicted set, thereby guaranteeing $(1+\epsilon)$-consistency. On the other hand, if some subset of the predicted set is served, Lemma 3 guarantees that the total social welfare obtained by Mechanism 2 is within a factor $\frac{9}{100(2\gamma+1)H_{n}}$ of the optimal social welfare. Since $H_{n}=O(\log{n})$ and $\gamma$ is a constant for any choice of constant $\epsilon$ we obtain $O(\log{n})$-robustness. Finally, if some subset of an unpredicted set is served, Lemma 5 guarantees that Mechanism 2 obtains welfare within a factor $\frac{9}{10(2\gamma+1)H_{n}}+\frac{1}{2H_{n}}$ of the optimal welfare. Again, since $H_{n}=O(\log{n})$ and $\gamma$ is constant for any constant $\epsilon$ we obtain $O(\log{n})$-robustness in this case. ∎
While the focus of our work is on leveraging advice to circumvent the informational limitations faced by clock auctions (recall that the lower bound from [DGR17] precludes clock auctions from achieving a $\log^{\tau}(n)$-approximation for any constant $\tau<1$), FollowTheUnpredictedLeader could still require exponential time for certain feasibility constraints (e.g., when $\mathcal{F}$ corresponds to the independent sets of a graph). Fortunately, a key advantage of FollowTheUnpredictedLeader is that it admits the use of a black-box approximation algorithm for the welfare and revenue maximization steps in lines 7 and 9. Specifically, if there exists an algorithm that outputs a set of rejected bidders which gives a $\rho$-approximation to the welfare of the best-rejected set in line 7 and an active set which gives $\rho$-approximation to the revenue in line 9, then our algorithm achieves robustness that scales linearly with $\rho$ (i.e., robustness becomes $O\left(\rho\cdot\frac{1}{\epsilon}\log n\right)$). This result is formalized in Corollary 1. For completeness, a full proof is provided in Appendix A.
Corollary 1.
For any downward-closed set system, Mechanism 2 with parameter $\gamma=\frac{10(1+\epsilon)}{9\epsilon}$ equipped with a $\rho$-approximate algorithm for the underlying problem of maximizing revenue/welfare (at some fixed clock prices) is $(1+\epsilon)$-consistent and $O(\rho\cdot\frac{1}{\epsilon}\cdot\log{n})$-robust for any constant $\epsilon>0$.